# CUB: Benchmarking Context Utilisation Techniques for Language Models

### v1

@@ Abstract
Incorporating external knowledge is crucial for knowledge-intensive tasks, such as question answering and fact checking. However, language models (LMs) may ignore relevant information that contradicts outdated parametric memory or be distracted by irrelevant contexts. While many context utilisation manipulation techniques (CMTs) that encourage or suppress context utilisation have recently been proposed to alleviate these issues, few have seen systematic comparison. In this paper, we develop CUB (Context Utilisation Benchmark) to help practitioners within retrieval-augmented generation (RAG) identify the best CMT for their needs. CUB allows for rigorous testing on three distinct context types, observed to capture key challenges in realistic context utilisation scenarios. With this benchmark, we evaluate seven state-of-the-art methods, representative of the main categories of CMTs, across three diverse datasets and tasks, applied to nine LMs. Our results show that most of the existing CMTs struggle to handle the full set of types of contexts that may be encountered in real-world retrieval-augmented scenarios. Moreover, we find that many CMTs display an inflated performance on simple synthesised datasets, compared to more realistic datasets with naturally occurring samples. Altogether, our results show the need for holistic tests of CMTs and the development of CMTs that can handle multiple context types.
@@ 1 Introduction
As a consequence, many different methods for increasing or suppressing LM context utilisation, henceforth referred to as CMTs (Context utilisation Manipulation Techniques), have been proposed. The methods encompass a broad range of approaches, from different decoding methods (Shi et al., 2024; Kim et al., 2024) to fine-tuning methods (Li et al., 2023), prompting (Liu et al., 2023), multi-agent (Feng et al., 2024; Du et al., 2024), and mechanistic interventions (Ortu et al., 2024; Jin et al., 2024). While each method yields promising results in isolation, their evaluation is often limited to narrow or idealised settings, leaving open the question of which approaches are applicable in real-world RAG scenarios. To address this evaluation gap, we develop a comprehensive CMT benchmark to test and compare different CMTs on datasets representative of different domains and tasks (Figure 1). Our contributions are as follows:
  • We develop CUB (Context Utilisation Benchmark) to allow for a comprehensive evaluation and comparison of CMTs (§3). [^1] Code will be available upon publication. CUB systematically tests the sensitivity of CMTs to underlying model and naturally occurring context types (gold, conflicting and irrelevant) on tasks representative of synthesised and realistic RAG scenarios.
  • We evaluate a cohort of state-of-the-art CMTs representative of the main categories of CMTs (§4) on our benchmark (§6).
  • We provide a deeper analysis of what CMT works best for a given scenario and identify areas of improvement for CMTs. We find that CMTs struggle to optimise performance across all context types, e.g. one approach may improve robustness to irrelevant contexts but degrade the utilisation of relevant contexts. This points to the need of CMTs that work well across all context types.
@@ 2 Related Work
Previous work has typically evaluated different CMTs on either of the dataset categories. CUB incorporates datasets representative of both knowledge-intensive tasks and diagnostic datasets, thus enabling comprehensive evaluations of CMTs in different settings.
@@
Many context utilisation manipulation techniques have recently been proposed. Existing CMTs can be categorised into one of four main groups based on intervention level, i.e. what aspect of the model they manipulate. 1) fine-tuning CMTs update model parameters to modify context utilisation. For example, fine-tuning on distracting contexts was found to yield improved robustness to distracting contexts (Li et al., 2023; Shen et al., 2024; Yoran et al., 2024). Moreover, Fang et al. (2024) specifically focus on different types of retrieval noise likely to be encountered in real-world environments and develop a fine-tuning approach to handle these. 2) prompting techniques modify the input to the LM to improve context utilisation, representing minimally modified settings. 3) mechanistic interventions on the LM modify certain model components at inference time to alter context utilisation. Examples involve attention modification (Ortu et al., 2024; Jin et al., 2024) and SpARe interventions (Zhao et al., 2025). Lastly, 4) decoding methods involve a modified decoding approach, applied to the output logits, to manipulate context utilisation. Examples include context-aware contrastive decoding (Yuan et al., 2024; Kim et al., 2024; Shi et al., 2024; Wang et al., 2024; Zhao et al., 2024) and lookback lens decoding (Chuang et al., 2024).
Apart from intervention level, many of the CMTs have different objectives, focused on improving one or multiple aspects of context utilisation. CMTs may focus on improving robustness to irrelevant contexts, faithfulness to conflicting contexts, or faithfulness to contexts in general.
Previous work has mainly focused on evaluating one CMT at a time, potentially due to the lack of a unified benchmark for CMTs. In this paper, we evaluate representatives from each of the four main categories of CMTs on CUB, comparing a total of seven CMTs.
@@
To the knowledge of the authors, there is not yet a benchmark for CMTs. The closest examples of existing benchmarks are RAG-Bench by Fang et al. (2024), KILT by Petroni et al. (2021) and AxBench by Wu et al. (2025). The first evaluates the retrieval-noise robustness of LMs, the second performance of RAG systems as a whole, and the latter steering techniques for LMs, focusing on safety and reliability. CUB takes inspiration from these benchmarks to create a comprehensive and relevant benchmark for the evaluation of CMTs.
@@ 3 CUB: A Context Utilisation Benchmark
Table 1: Statistics of the datasets that form CUB. ‘Conflict.’ denotes conflicting contexts and ‘Irrel.’ irrelevant contexts.
Given a CMT, CUB is designed to test the technique across different datasets, models and metrics. To unify the tests, CUB also incorporates a pre-defined method for the hyperparameter search of the CMT.
3.1 Language Models
CUB evaluates the model sensitivity of CMTs on up to nine different LMs. The open-sourced models covered by the benchmark are GPT-2 XL, Pythia (6.9B), Qwen2.5 1.5B, Qwen2.5 7B, Qwen2.5 32B (Radford et al., 2019; Biderman et al., 2023; Yang et al., 2024). For the Qwen models we include the instruction-tuned variants. We also evaluate the API-based LLM Cohere Command A with  [^1] B parameters. [^2] https://cohere.com/blog/command-a The model selection is performed to enable comparisons across model families, model sizes, instruction-tuning and API-based LLMs. However, all LMs are not compatible with all CMTs evaluated on CUB – the selection of LMs onto which a CMT is applied depends on the CMT, further explained in Section 4. In addition, we adapt the prompts in CUB with prompt templates compatible with each model type under consideration (base, instruction-tuned and chat-API).
@@ 3.2 Datasets
To evaluate how CMTs respond to different types of contextual information, CUB evaluates each CMT on CounterFact, NQ and DRUID (see Table 1). [^3] The datasets can be found on Hugging Face Datasets under copenlu/{cub-counterfact,cub-nq,cub-druid}. The inclusion of these datasets is based on three key criteria: (i) diversity in task difficulty, (ii) diversity in realistic and synthesised RAG scenarios, and (iii) high utilisation in related work. CounterFact represents a causal language modelling task based on a controlled setup with simple counterfactual contexts synthesised to conflict with model memory. NQ represents a popular, and more realistic setup, focused on RAG for open-domain QA of greater difficulty with contexts sampled from Wikipedia. DRUID is a fairly new dataset, representing another important RAG task – that of automated fact-checking; this requires a greater level of reasoning based on naturally occurring claims and evidence sampled from the internet. While DRUID has yet to see widespread use in studies of context utilisation, we include it in CUB as it is one of few datasets closely aligned with real-world RAG scenarios.
For each dataset, we curate samples representative of the three types of contexts that may be encountered in realistic RAG scenarios: 1) gold contexts that are relevant and do not contradict LM memory, 2) conflicting contexts that are relevant but contradict LM memory or gold labels, and 3) irrelevant contexts that should be ignored by the LM (Fang et al., 2024). For each dataset, we sample validation and test splits. To allow for fair and unified comparisons between CMTs, the validation set is used to tune potential hyperparameters of the CMT under evaluation. The test split is used for the final evaluation. More details on the datasets can be found in Appendix B.
@@
To construct a CounterFact dataset with counterfactual contexts, we first identify samples from LAMA that have been memorised by Pythia 6.9B, following the approach by Saynova et al. (2025). We base the CounterFact dataset on Pythia to obtain a set of samples likely to have been memorised by all CUB models, since LMs have been found to memorise more facts as they grow in size (Saynova et al., 2025). We confirm this in Appendix B; all CUB LMs are found to have memorised at least 70% of the CounterFact samples. Based on the known fact triplets, we sample conflicting contexts following the approach of Meng et al. (2022). We also sample gold contexts that simply state the correct triplet. For the irrelevant contexts, we randomly sample fact triplets unrelated to the sample query.
@@
The gold context samples are simply the original NQ samples. For the collection of samples with conflicting contexts, we follow a substitution approach inspired by the method of Longpre et al. (2021). We create conflicting contexts that promote a different answer simply by taking the gold context and substituting the gold answer in the context. The substitute answer is sampled to yield coherent conflicting contexts, and to have a different meaning compared to the gold answer. For the collection of samples with irrelevant contexts, we apply a LM re-ranker to identify the most relevant non-gold paragraph from the Wikipedia page in which the gold context was found. With this approach, we collect irrelevant contexts representative of real-world RAG scenarios.
@@
The <claim, evidence> samples of DRUID have been manually annotated for stance of the evidence (supports, refutes, insufficient or irrelevant). We map stance to context type as described in Appendix B. No context synthesis is necessary for the DRUID samples as they, by virtue of utilising naturally occurring samples from a RAG pipeline, already contain samples representative of gold, conflicting and irrelevant contexts. Moreover, since DRUID represents a reasoning task, asking the model whether provided evidence supports the claim under consideration (True or False), or is insufficient (None), the output space for the DRUID samples is limited to three tokens (True, False or None).
@@ 3.3 Metrics
Similarly to Jin et al. (2024) we use a binary score to measure context utilisation. We refer to it as the binary context utilisation ($\mathrm{BCU}$) score and define it as follows. For relevant contexts (gold and conflicting) the score is 1 if the LM prediction is the same as the token promoted by the context, $t_{C}$, and 0 otherwise. For irrelevant contexts the score is 1 if the LM prediction is the same as the memory token, $t_{M}$, (i.e. the prediction made by the model before any context has been introduced) and 0 otherwise. We report the averaged $\mathrm{BCU}$ score per context type. To assess the relative effectiveness of CMTs, we also report the net gain of each CMT, compared to when no CMT is applied, using $\mathrm{BCU}$ score ($\Delta=\mathrm{BCU_{\texttt{CMT}}}-\mathrm{BCU_{\texttt{Regular}}}$). We also consider continuous context utilisation, $\mathrm{CCU}$, a more fine-grained metric that measures the change in outputted token probabilities as context is introduced. Appendix C contains more details on the metric.
We also measure the accuracy of each method. For CounterFact and DRUID, accuracy is measured based on whether the first generated token is the same as the first gold token. For NQ, for which the correct answer may be different permutations of the same set of tokens, we measure accuracy based on whether the first output token (e.g. “July”) matches any of the tokens in the answer (e.g. “15 July”).
@@ 3.4 Hyperparameter Search
For CMTs requiring hyperparameter tuning, we use the validation set of each dataset to select values that maximise the average BCU across all context types, unless a method-specific tuning procedure is explicitly specified. This ensures a fair comparison between CMTs. Further details are shown in Appendix D.
@@ 4 Context Utilisation Manipulation Techniques
We benchmark a total of seven different CMTs on CUB, all of which are state-of-the-art representatives from the main categories of CMTs. Table 2 summarises the key characteristics of the CMTs, including their main objective, intervention level, and cost in terms of tuning and inference. As a baseline, we also evaluate regular LMs on the same input, with no CMT applied (Regular).
@@
We adapt the approach of Li et al. (2023), which fine-tunes LMs to ensure the usage of relevant contexts. It considers four different types of contexts: relevant, irrelevant, empty, and counterfactual contexts. To align the domain with our evaluation data, we curate the fine-tuning data with two QA datasets Joshi et al. (2017); Rajpurkar et al. (2018), one FC dataset Schlichtkrull et al. (2023), and one sentence completion dataset Marjanovic et al. (2024). Before fine-tuning each LM, we elicit its parametric answers by querying without contexts. We then select the questions that the LM answered correctly and pair them with irrelevant and empty contexts. The fine-tuning data thus contains contexts that can be irrelevant, counterfactual, or empty. During fine-tuning, we train the LM to generate answers aligned with the provided context. When the context is irrelevant, we train the LM to be robust, i.e. ignore the context and output its parametric answer. Due to the computational costs associated with fine-tuning billion-sized LMs, we use the Low-Rank Adaptation method Hu et al. (2021). Additional details can be found in Appendix E.
@@
We curate a set of 12 prompts for each evaluation dataset and optimise the prompt selection to each evaluated model. Each set of prompts is based on 6 prompts curated by a human, similarly to the approach by Jin et al. (2024), and 6 prompts generated by a LLM, [^4] Mainly by ChatGPT, but also by Microsoft Co-pilot. similarly to the approach by Wu et al. (2025).
@@
Figure 2: Overview of the multi-agent approach.
Inspired by LM agents and self-refinement (Du et al., 2024; Feng et al., 2024; Madaan et al., 2023), which are widely adopted techniques in reasoning tasks, we decompose context utilisation into two components – relevance and context faithfulness – and assign each as a separate task to an individual LM agent. We aim to examine whether LMs are capable of accurately evaluating context relevance and answer faithfulness, to subsequently self-correct themselves for improved faithfulness to relevant contexts. As illustrated in Figure 2, we first assess relevance using the relevance agent to determine whether the provided context should be used. Then, the faithfulness agent provides feedback on the model response that was generated with context. If the feedback indicates that the initial answer is unfaithful, the model generates a self-refined answer based on that feedback. Given that these tasks require instruction-following capabilities, we restrict our evaluation to instruction-tuned or chat LMs. Further details can be found in Appendix F.
@@
We adopt the PH3 method by Jin et al. (2024). The method is implemented in two steps: 1) identification of attention heads responsible for context or memory reliance via path patching and 2) pruning the identified attention heads for increased memory or context usage. To identify attention heads, we use the CounterFact datasets with samples that elicit exact fact recall in each studied model (Saynova et al., 2025). For the evaluation on our studied datasets, we tune the number of heads to prune on the validation splits of each evaluation dataset, similarly to the approach by Jin et al. (2024). PH3 can be used in two different modes – suppressing context attention heads or suppressing memory attention heads. We tune the attention head configuration for each mode and report the results (PH3 +context enhances context utilisation by the suppression of memory heads, and vice versa for PH3 +memory).

### v2

@@ Abstract
Incorporating external knowledge is crucial for knowledge-intensive tasks, such as question answering and fact checking. However, language models (LMs) may ignore relevant information that contradicts outdated parametric memory or be distracted by irrelevant contexts. While many context utilisation manipulation techniques (CMTs) have recently been proposed to alleviate these issues, few have seen systematic comparison. In this paper, we develop CUB (Context Utilisation Benchmark) – the first comprehensive benchmark designed to help practitioners within retrieval-augmented generation (RAG) diagnose CMTs under different context conditions. With this benchmark, we conduct the most extensive evaluation to date of seven state-of-the-art methods, representative of the main categories of CMTs, across three diverse datasets and tasks, applied to nine LMs. Our results reveal that most existing CMTs struggle to handle the full spectrum of context types encountered in real-world retrieval-augmented scenarios. We also find that many CMTs display inflated performance on simple synthesised datasets, compared to more realistic datasets with naturally occurring samples. Our findings expose critical gaps in current CMT evaluation practices and demonstrate the need for holistic testing and the development of CMTs that can robustly handle multiple context types.
@@ 1 Introduction
As a consequence, many different methods for increasing or suppressing LM context utilisation, henceforth referred to as CMTs (Context utilisation Manipulation Techniques), have been proposed. The methods encompass a broad range of approaches, from different decoding methods (Shi et al., 2024; Kim et al., 2024) to fine-tuning methods (Li et al., 2023), prompting (Liu et al., 2023), multi-agent (Feng et al., 2024; Du et al., 2024), and mechanistic interventions (Ortu et al., 2024; Jin et al., 2024). Many of these CMTs are developed with distinct objectives in mind, often targeting a specific aspect of context utilisation – such as improving robustness to irrelevant contexts or enhancing faithfulness to conflicting information. While each method yields promising results in isolation, their evaluation is often limited to settings that closely match the method’s target objective, making it unclear whether these approaches will generalise or remain effective in real-world RAG scenarios. To address this pressing evaluation gap, we develop the first comprehensive CMT benchmark to systematically test and compare different CMTs across datasets that truly represent the diversity of real-world domains and tasks (Figure˜1). Our contributions are as follows:
  • We develop CUB (Context Utilisation Benchmark) – the first comprehensive benchmark designed to enable rigorous evaluation and comparison of CMTs under various context conditions (§3). [^1] Code will be made available upon publication. CUB evaluates CMTs across gold, conflicting, and irrelevant contexts that capture key RAG challenges.
  • We conduct the most extensive evaluation to date of CMTs, assessing seven state-of-the-art methods representative of the main categories of CMTs (§4) across our benchmark (§6). The study encompasses approximately 750 experimental data points, spanning 9 LMs, 7 CMTs, 3 datasets, and 3 different context types.
  • We provide insights into what CMT works best for different scenarios and identify critical areas for improvement. Our analysis reveals that existing CMTs face fundamental trade-offs – they struggle to optimise performance across all context types. This points to a need for next-generation CMTs that can robustly handle diverse context conditions.
@@ 2 Related Work
Previous work has typically evaluated different CMTs on either of the dataset categories in isolation, creating a fragmented understanding of CMT effectiveness. CUB bridges this critical gap by incorporating datasets representative of both knowledge-intensive tasks and diagnostic datasets, enabling the first comprehensive evaluation of CMTs across truly diverse settings.
@@
Many context utilisation manipulation techniques have recently been proposed. Existing CMTs can be categorised into one of four main groups based on intervention level, i.e. what aspect of the model they manipulate. 1) fine-tuning CMTs update model parameters to modify context utilisation. For example, fine-tuning on distracting contexts was found to yield improved robustness to distracting contexts (Li et al., 2023; Shen et al., 2024; Yoran et al., 2024). Fang et al. (2024) specifically focus on different types of retrieval noise likely to be encountered in real-world environments and develop a fine-tuning approach to handle these. 2) prompting techniques modify the input to the LM to improve context utilisation, representing minimally modified settings. 3) mechanistic interventions on the LM modify certain model components at inference time to alter context utilisation. Examples involve attention modification (Ortu et al., 2024; Jin et al., 2024) and SpARe interventions (Zhao et al., 2025). Lastly, 4) decoding methods involve a modified decoding approach, applied to the output logits, to manipulate context utilisation. Examples include context-aware contrastive decoding (Yuan et al., 2024; Kim et al., 2024; Shi et al., 2024; Wang et al., 2024; Zhao et al., 2024) and lookback lens decoding (Chuang et al., 2024).
Despite significant progress in developing new CMTs, systematic evaluations remain scarce. Prior studies have largely assessed individual CMTs in isolation, likely due to the lack of a unified benchmark. In this paper, we address this gap by conducting the first comprehensive evaluation of seven representative CMTs across the four main categories of CMTs, using the CUB dataset.
@@
To our knowledge, there is not yet a benchmark specifically designed for CMTs, representing a significant research gap. The closest examples of existing benchmarks are RAG-Bench by Fang et al. (2024), KILT by Petroni et al. (2021) and AxBench by Wu et al. (2025). However, these serve different purposes: the first evaluates the retrieval-noise robustness of LMs, the second assesses performance of RAG systems as a whole, and the latter focuses on steering techniques for LMs, emphasising safety and reliability rather than context utilisation. CUB addresses this critical infrastructure gap by creating the first comprehensive and purpose-built benchmark specifically for the evaluation of CMTs, taking inspiration from these existing benchmarks while addressing the unique challenges of CMT evaluation.
@@ 3 CUB: A Context Utilisation Benchmark
Given a CMT, CUB is designed to systematically test the technique across different datasets, models and metrics, providing rigorous evaluation. To ensure fair and meaningful comparisons, CUB incorporates a carefully designed pre-defined method for hyperparameter search, eliminating potential bias in method comparisons.
3.1 Language Models
CUB evaluates the model sensitivity of CMTs across a range of nine different LMs, providing insights into how CMT effectiveness varies with model architecture and scale. The comprehensive model coverage includes both open-sourced models – GPT-2 XL, Pythia (6.9B), Qwen2.5 (1.5B, 7B, and 32B) (Radford et al., 2019; Biderman et al., 2023; Yang et al., 2024) – and API-based Cohere Command A ( [^1] B), which is specialised for RAG. [^2] https://cohere.com/blog/command-a For the Qwen models, we strategically include both base and instruction-tuned variants to capture the impact of instruction tuning on context utilisation. This carefully curated model selection enables comprehensive comparisons across model families, model sizes, instruction-tuning approaches, and deployment paradigms.
@@ 3.1 Language Models
Table 1: Statistics of the datasets that form CUB. ‘Conflict.’ and ‘Irrel.’ denote conflicting and irrelevant contexts, respectively.
@@ 3.2 Datasets
To systematically evaluate how CMTs respond to different types of contextual information, CUB evaluates each CMT across CounterFact, NQ and DRUID (see Table˜1). [^3] The datasets can be found on Hugging Face Datasets under copenlu/{cub-counterfact,cub-nq,cub-druid}. The inclusion of these datasets is based on three key criteria that ensure comprehensive coverage: (i) diversity in task difficulty, (ii) diversity in realistic versus synthesised RAG scenarios, and (iii) high utilisation in related work, ensuring our findings connect to the broader literature. CounterFact represents a controlled causal language modelling task with carefully synthesised counterfactual contexts designed to conflict with model memory. NQ represents a popular and more realistic setup focused on RAG for open-domain QA of substantially greater difficulty, with contexts sampled from Wikipedia. DRUID is a cutting-edge dataset representing another critical RAG task – automated fact-checking – which requires sophisticated reasoning based on naturally occurring claims and evidence from the internet.
For each dataset, we curate samples representative of the three types of contexts that are fundamental to realistic RAG scenarios: 1) gold contexts that are relevant, 2) conflicting contexts that are relevant but always contradict the LM’s memory, and 3) irrelevant contexts that provide no information to solve the given question (Fang et al., 2024). This approach ensures that CUB captures the full spectrum of challenges to RAG in real-world deployments. More details on the data collection process can be found below and in Appendix˜B.
@@
To construct a CounterFact dataset with counterfactual contexts, we first identify samples from LAMA that have been memorised by Pythia 6.9B, following the approach by Saynova et al. (2025). We base the CounterFact dataset on Pythia to obtain a set of samples likely to have been memorised by all CUB models, since LMs have been found to memorise more facts as they grow in size (Saynova et al., 2025). We confirm this in Appendix˜B; all CUB LMs are found to have memorised at least 70% of the CounterFact samples. Based on the known fact triplets, we sample conflicting contexts following the approach of Meng et al. (2022). We also sample gold contexts that simply state the correct triplet. For the irrelevant contexts, we randomly sample fact triplets unrelated to the sample query.
@@
The gold context samples are simply the original NQ samples. For the collection of samples with conflicting contexts, we follow a substitution approach inspired by the method of Longpre et al. (2021). For the collection of samples with irrelevant contexts, we apply a LM re-ranker to identify the most relevant non-gold paragraph from the Wikipedia page in which the gold context was found. With this approach, we collect irrelevant adversarial contexts representative of real-world RAG scenarios.
@@
[…]
@@ 3.3 Metrics
Similarly to Jin et al. (2024) we use a binary score to measure context utilisation. We refer to it as the binary context utilisation ($\mathrm{BCU}$) score and define it as follows. For relevant contexts (gold and conflicting) the score is 1 if the LM prediction is the same as the token promoted by the context, $t_{C}$, and 0 otherwise. For irrelevant contexts the score is 1 if the LM prediction is the same as the memory token, $t_{M}$, (i.e. the prediction made by the model before any context has been introduced) and 0 otherwise. We report the averaged $\mathrm{BCU}$ score per context type. To assess the relative effectiveness of CMTs, we also report the net gain of each CMT, compared to when no CMT is applied, using $\mathrm{BCU}$ score ($\Delta=\mathrm{BCU_{\texttt{CMT}}}-\mathrm{BCU_{\texttt{Regular}}}$). We also consider continuous context utilisation, $\mathrm{CCU}$, a more fine-grained metric that measures the change in outputted token probabilities as context is introduced, further described in Appendix˜C.
@@ 3.4 Hyperparameter Search
For CMTs requiring hyperparameter tuning, we use the validation set of each dataset to select values that maximise the average BCU across all context types, unless a method-specific tuning procedure is explicitly specified (see Appendix˜D). This ensures a fair comparison between CMTs.
@@ 4 Context Utilisation Manipulation Techniques
We benchmark a comprehensive set of seven state-of-the-art CMTs on CUB, strategically selected to represent all major categories of CMTs and provide unprecedented coverage of the CMT landscape. Table 2 provides a detailed overview of the key characteristics of each CMT, including their primary objectives, intervention levels, and computational costs for both tuning and inference. We adaptively select the appropriate set of LMs for each CMT based on technical compatibility. As a baseline for comparison, we also evaluate regular LMs on identical inputs with no CMT applied (Regular), enabling direct assessment of CMT effectiveness.
@@
We adapt the approach of Li et al. (2023), which fine-tunes LMs to ensure the usage of relevant contexts (see Appendix˜E). It considers four different types of contexts: relevant, irrelevant, empty, and conflicting contexts.
@@
We curate a set of 12 prompts for each evaluation dataset, tailoring prompt selection to each evaluated model’s characteristics. Each prompt set combines human expertise and AI generation: 6 prompts are curated by human experts following established methodologies (Jin et al., 2024), while 6 prompts are generated by advanced LLMs, [^4] Mainly by ChatGPT, but also by Microsoft Co-pilot. following established prompt engineering approaches (Wu et al., 2025).
@@
Inspired by LM agents and self-refinement (Du et al., 2024; Feng et al., 2024; Madaan et al., 2023), which are widely adopted techniques in reasoning tasks, we decompose context utilisation into two components – relevance and context faithfulness – and assign each as a separate task to an individual LM agent (see Appendix˜F). We aim to examine whether LMs are capable of accurately evaluating context relevance and answer faithfulness, to subsequently self-correct themselves for improved faithfulness to relevant contexts.
@@
We adopt the PH3 method by Jin et al. (2024) (see Appendix˜G). PH3 can be used in two different modes – suppressing context attention heads (PH3 +memory) or suppressing memory attention heads (PH3 +context).